use elimination to find all points of intersection of the graphs of 3x+2y=10 and 2x+5y=3
Answer by sofiyac(983)

3x+2y=10 multiply this one by -2
2x+5y=3 multiply this one by 3, then add the two new equations
-
-6x-4y=-20
6x+15y=9 now add them
-----------
11y=-11 divide each side by 11
y=-1 next plug that into either of the original equations and solve for x, i'm going to work with the first one
3x+2(-1)=10
3x-2=10 add 2 to each side
3x=12 divide each side by 3
x=4
(4,-1)